Solve the equations for x:
11(a+3b)−2(5a−5x)−4(3a+8x)=0;

Answer by mananth(16946)

11(a+3b)−2(5a−5x)−4(3a+8x)=0;

11a+33b-10a+10x-12a-32x=0
-11a=33b -22x=0
22x = -11a-33b
22x=-11(a+3b)
x = -11/22 (a+3b)
x= -1/2*(a+3b)